In a lottery game, a player picks six numbers from 1 to 27. If the player matches all six numbers, they win 50,000 dollars. Otherwise, they lose $1.

What is the expected value of this game? $
1
Expert's answer
2020-07-03T16:15:05-0400

Solution:

E[X]=x1p1+x2p2

x1=50000,

p1=m/n, m=1, n=binomial(27,6)=27!/6!/21!=296010,

here binomial(27,6) is the binomial coefficient (276).

p1=1/296010,

x2=-1,

p2=1-p1=1-1/296010=296009/296010.

E[X]=50000*1/296010+(-1)296009/296010=-82003/98670=-0.83...

Answer:

E[X]=-82003/98670=-0.83...